OpenStudy (anikate):

Find the center and radius. x^2 + y^2 +20x -2y -92=0

OpenStudy (anonymous):

you need to apply completing the square here, for both x and y, then you can establish the normal equation of a circle.

OpenStudy (anonymous):

\[x^2 + 20x + x^2 - 2y = 92\] \[x^2 + 20x + (10)^2 + y^2 - 2y + (1)^2 = 92 + (10)^2 + (1)^2\] \[\large (x + 10)^2 + (y - 1)^2 = (\sqrt{193})^2\] Here radius is \(\sqrt{193} \approx 13.9\) Center is : \((-10, 1)\)

OpenStudy (anikate):

why is there a sq root on top of the 193? @waterineyes

OpenStudy (anikate):

@karatechopper

OpenStudy (anikate):

@GOODMAN

OpenStudy (anikate):

hey can u plz explain to me why there is a sq rt over the 193 and the sqaured

OpenStudy (anonymous):

Because you need to find the radius. And the radius is the diameter squared. So 193 was the diameter, but we need the radius. So @waterineyes squared it..

OpenStudy (anikate):

ok

OpenStudy (anikate):

@Calcmathlete can you plz explain this same question again without bothering about water in eye's posts

OpenStudy (anikate):

@saifoo.khan can u plz explain this question

OpenStudy (saifoo.khan):

Sorry, i never studied eclipse. :/

OpenStudy (anikate):

huh? how but ur so smart?

OpenStudy (anikate):

calc plz help me

OpenStudy (anonymous):

lol @saifoo.khan This is a circle XD Well, I guess a circle is a type of ellipse

OpenStudy (anonymous):

Alright. Hold on.

OpenStudy (anonymous):

Do you understand what water above did? She completed the square. Are you familiar with it?

OpenStudy (anikate):

yep with completing the square

OpenStudy (saifoo.khan):

lol. Yup! similar things..

OpenStudy (anonymous):

Do yuo understand how it works though?

OpenStudy (anikate):

nope I think she messed up somewhere

OpenStudy (anonymous):

Ok. Let's see... \[x^2 + y^2 + 20x - 2y - 92 = 0\]\[x^2 + y^2 + 20x - 2y = 92\]\[x^2 + 20x + y^2 - 2y = 92\]\[x^2 + 20x + 100 + y^2 - 2y + 1 = 92 + 100 + 1\]\[(x + 10)^2 + (y - 1)^2 = 193\]THis is the form that you need it in. The form for a circle is \[(x - h)^2 + (y - k)^2 = r^2\]Do yuo see the resemblance?

OpenStudy (anonymous):

The center of a circle is (-h, -k) and the radius is 4. Do you know how to get them now?

OpenStudy (anonymous):

I mean the radius is r.

OpenStudy (anonymous):

Are you following???

OpenStudy (anikate):

alright sry, my mom called me

OpenStudy (anikate):

sryy to keep you waiting

OpenStudy (anonymous):

Do you get what to do from there?

OpenStudy (anikate):

yes I already go the answer

OpenStudy (anikate):

thx though

OpenStudy (anonymous):

np :)